A rectangle is 6cm long and 5 cm wide. When each dimension is increased by x cm, the area is tripled. Find the value of x.
Found 2 solutions by stanbon, Greenfinch:
Answer by stanbon(75887)

A rectangle is 6cm long and 5 cm wide. When each dimension is increased by x cm, the area is tripled. Find the value of x.
------------------------------
Original area = 6*5 = 30 cm^2
---------------------------------------
Equation:
(6+x)(5+x) = 3*30
x^2 + 11x + 30 - 90 = 0
x^2 + 11x - 60 = 0
-------------
(x+15)(x-4) = 0
Positive solution:
x = 4 cm
=============================
Cheers,
Stan H.

Answer by Greenfinch(383)

Initially 6 x 5 = 30 sq cm
In the new case (x + 6)(x + 5)=90
x^2 + 11x + 30 = 90 or
x^2 + 11x - 60 = 0
(x - 4)(x + 15) = 0
x = 4 or -15
The number must be positive, so x=4
this gives 10 and 9 cm which give an area of 90 cm sq